36% of women consider themselves fans of professional basebal You random y select six women and ask each if she considers herself a fan of professional baseball Complete pats a thou h d below (a) Find the mean of the binomial distribution. (Round to the nearest tenth as needed) (Round to the nearest tenth as needed.) (b) Find the variance of the binomial distribution. 2(Round to the nearest tenth as needed.) (c) Find the standard devistion of the binomial distribution. -(Round to the nearest tenth as needed,) (d) Interpret the results in the context of the real-life situation. On average. out of 6 women would consider themselves basebal fans. The standard deviation is women, so in most samples of 6 women, the number of women who consider themselves baseball ans would differ from the mean by no more than Type integers or decimals rounded to the nearest tenth as needed.)
Answer #1

Here we have

n=6, p=0.36

(a)

The mean is

μ np = 6 . 0.36 2.16 ~ 2.2

(b)

The variance is

σ2 = np(1-p) = 6 . 0.36 . 0.64 ~ 1.4

(c)

The standard deviation is

б . 0.50 , 0,б

(d)

On average, 2.2 out of 6 women would consider themselves baseball fans. The standard deviation is 1.2 women, so in most samples of 6 women, the number of women who consider themselves baseball fans would differ from the mean by no more than 2*1.2 = 2.4.